OpenStudy (anonymous):

Bill drove to the mountains last weekend. There was heavy traffic on the way there, and the trip took 6 hours. When Bill drove home, there was no traffic and the trip only took 4 hours. If his average rate was 22 miles per hour faster on the trip home, how far away does Bill live from the mountains?

OpenStudy (anonymous):

r t = d Let x and y be the speed up to the mountains and the speed back home respectively. Solve the following for x and y. {6 x = 4 y, x = y - 22} {x = 44, y = 66} 6*44 = 264 miles from the mountains.